1/6 - x/2 = x/3 - 5/3
Answer by Maths68(1474)

1/6 - x/2 = x/3 - 5/3
Multiply by 6 both sides
1-3x=2x-10
-3x-2x=-10-1
-5x=-11
5x=11
X=11/5
X=2 1/5